OpenStudy (anonymous):

Convert each equation to Vertex Formula and Graph each equation showing its vertex. x2−10x−11=0

OpenStudy (anonymous):

You need to complete the square first. x^2 - 10x - 11 = 0 x^2 - 10x + 25 - 11 = 0 + 25 (x - 5)^2 -36 = 0 Vertex (5,-36)